What is the standard form of the equation of the line with slope -4/7 passing through the point (-5,2)
Answer by sachi(548)

standard form of the equation of the line with slope -4/7 passing through the point (-5,2) is
[y-2]/[x-(-5)]=-4/7
or [y-2]/[x+5]=-4/7
or [y-2]=-4/7=-4x/7-20/7
or y=-4x/7-(20/7)+2=-4x/7-6/7
ans